# Elaborating Higgs to dimuon decay from gluon fusion by decorrelation and jet substructure

### v1

@@ Abstract
Discovery of the Higgs boson decay to dimuon is anticipated soon based on the current evidence. Precise categorization of the events without affecting the invariant mass shape is crucial in the analysis. Decorrelation of the invariant mass and the output of discriminators (the score of discriminators) is essential for consistent and precise analysis. In this paper we use distance correlation as the additional loss function to achieve the decorrelation for discriminators and examine various analysis methods. The analyses with and without jet substructure variables are presented. Adding jet substructure variables considerably improves the significance of the Higgs to dimuon signal from gluon fusion.
@@ I Introduction
After Higgs discovery Chatrchyan et al. (2012)Aad et al. (2012), various properties of Higgs have been measured including Higgs couplings to Standard Model (SM) particles. The SM predicts the Higgs coupling to a particle is proportional to its mass. Accordingly, Higgs couplings to heavy particles, such as gauge bosons and third generation fermions, are well measured Aad et al. (2020)Aaboud et al. (2018)Sirunyan et al. (2018)Aaboud et al. (2019). However, couplings to relatively light particles have not been satisfactorily measured so far. The Higgs to dimuon channel is the most sensitive channel among the second generation fermions at the LHC due to the relatively clean leptonic final states. The property of Higgs couplings to light fermions can verify the Higgs mechanism of the SM more accurately. Furthermore, the growing interest towards a muon collider makes this channel more interesting.
@@
A crucial requirement for the analyses using categorization above is that the information about the invariant mass should not be used in the categorization process. In other words, the output of discriminators, or the discriminator score, should be statistically independent of the invariant mass. Categorization by a score which is statistically dependent on the invariant mass generally results in different invariant mass distributions for each category, $f(m|\text{score})\neq f(m)$, where $f$ represents the probability density function. If the invariant mass distribution is distorted, the number of events extracted from each category can not be reliably trusted. In extreme cases, a resonance peak near the Higgs mass may appear in the background $m_{\mu\mu}$ distribution (See Fig.1 for example) adding a significant uncertainties on the number of signal events near the Higgs bump region. In addition, if we fit to the invariant mass after the categorization by a discriminator correlated to the invariant mass, it results in double usage of the invariant mass, compounding the error. Therefore the standard likelihood fitting to the invariant mass distribution can be consistently combined with machine discriminator results only when two results are decorrelated/deassociated.  [^1] Technically, the term ‘deassociate/deassociation’ is more appropriate Altman and Krzywinski (2015), but we will use the term ‘decorrelate/decorrelation’ following other literature Kasieczka and Shih (2020)
@@
To avoid the distortions in the invariant mass distribution that arise when events are selected by a discriminator correlated with the invariant mass, researchers have not only securely excluded the invariant mass but also carefully selected other variables to make the categories as insensitive to the invariant mass as possible. However, as will be seen in the paper, machine discriminators generally learn about $m_{\mu\mu}$ since it has strong discrimination power. An ad hoc choice of variables which seemingly unrelated to the $m_{\mu\mu}$, qualitative check of distortion in distributions after categorization Bourilkov et al. (2019) and a posteriori modification of fitting functions are not satisfactory solutions. In this paper we provide a quantitative and consistent framework to deal with this problem. Among the various measures of dependence between two sets of feature variables, we choose ‘distance correlation’ Kasieczka and Shih (2020) Székely et al. (2007) since it is differentiable, so being compatible with deep learning frameworks, well defined in both population and sample spaces, and capable of capturing both linear and non-linear dependence. At the same time the range is well defined such that it is 0 when there is no correlation and is 1 when it is fully deterministic/dependent. By leveraging the flexiblility of deep neural network (DNN), we use the distance correlation between $m_{\mu\mu}$ and output of discriminator as an additional loss for training the discriminator. As a result, it has been found that almost perfect statistical independence (for both signal and background data classes simultaneously) between $m_{\mu\mu}$ and the discriminator output can be systematically achieved, enabling us to choose the required degrees of independence, which is crucial for accurate analysis, despite being in a trade-off relation to the performance of the discriminator. By carefully balancing this trade-off, this approach ensures reliable categorization without compromising the integrity of the invariant mass distribution.
@@
Compared to the analysis using only the kinematic variables of reconstructed particle objects, incorporating jeet substructure variables of the jets associated with the dimuon system as additional input features for the discriminator can provide extra power, especially in discriminating the gluon dominant ggH signal from other background processes. Currently the vector boson fusion (VBF) is the leading production channel for the Higgs to dimuon discovery despite the fact that the ggH production cross section is more than 10 times larger than VBF one at $\sqrt{s}=13$TeV CERN (2017). In Cho et al. (2020), it has been demonstrated that ggH channel can be exploited to improve the sensitivity of the Higgs invisible signal if jet substructure variables are fully utilized to distinguish the flavors and kinematics of ISR by DNN analysis. The difference in parton luminosities and QCD color factors of vertices results in the signal ggH ISR to be gluon-rich, while the background Drell-Yan (DY) ISR to be quark-rich. In addition, due to the helicity conservation, ggH ISR has a deficit of quark jets at the central rapidity region.  [^2] There is no issue of decorrelation in Higgs to invisible study, the treatment of systematic uncertainties originating from the simulation remain as open question. In this study, the same method is used to improve the sensitivity of Higgs to dimuon signal from ggH channel.
@@
The paper is composed as following. In section.II, we first outline the event simulation and ggH selection cuts. We also summarize the analysis strategy including training variables, common training schemes, and the categorization procedure. In section.III, we introduce three types of discriminators and compare their result in detail. In particular, with the $m_{\mu\mu}$ decorrelated discriminator implemented using the distance correlation, the degree of statistical independence and the discrimination performance will be compared. In section. IV, we compute the discovery significance using the categorized events and check the effect of including jet substructure variables.
@@ II Events and analysis strategy
When an event is classified being from ggH, it means that the event passed all selection cuts to identify ggH, regardless of genuine origin of the event. Therefore, all possible Higgs production channels have to be simulated. However, the contribution from other channels to the ggH signals is much smaller than genuine ggH channel due to their small cross sections. Next to leading order contribution is from VBF channel which is less than 5% of ggH channel. Thus we simulated only ggH and VBF channel. Similarly, there are many channels which act as backgrounds against ggH. Among them, we included the dominant background DY and $\bar{\text{t}}$ which has non-negligible contribution to multi-jet process (more than 20% of DY). All the events are simulated by MadGraph5_aMC@NLO Alwall et al. (2014) at $\sqrt{s}=13$TeV. Each simulation was done inclusively for the number of jets, so that we can properly include both soft and hard jet emissions.
After the simulation, the selection of events for ggH was conducted. Basically, an event is required to have two oppositely charged muons with $p_{T}>20$GeV and $|\eta|<2.4$. Any events including b-jet and/or additional charged lepton are vetoed to avoid overlap with $\bar{t}$ and VH channel. Finally, events with two energetic and largely separated jets ($m_{jj}>400$GeV, $|\Delta\eta_{jj}|>2.5)$ are rejected as those are to be identified as VBF.
@@
The discriminators are trained separately for events with different number of jets $n_{\text{jet}}$ because we have different number of training variables for each $n_{\text{jet}}$. The training sample consists of one million signals and backgrounds respectively, 70% for training and 30% for validation. We will train three types of machine discriminators to compare the results; BDT as a benchmark, (ordinary) DNN and decorrelated DNN. After the training, total data set is divided into categories according to the output of discriminators, or equivalently, by the signal efficiency. The boundaries are recursively selected to maximize the total expected significance with $S/\sqrt{B}$ as a figure of merit, while minimizing the number of categories (until the gain from additional boundary is less than 1%).
A typical next step is to extract the number of events from fitting of data to $m_{\mu\mu}$ distribution. However we will skip the extraction procedure and use the number of events by simulation directly because we do not have real observed data to fit. Since we already used large data, fitting to another set of simulated data is not worth much. Though our result depends on the simulated data, there is no significant problem in comparison between discrimination schemes and check the impact of including jet substructure variables. Therefore after the categorization, we will compute the discovery significance by profile likelihood fitting to the $m_{\mu\mu}$ distribution.
@@ III Discriminators
The astonishing success of DNN is based on the ability to build a hidden representation from the training variables. However, due to the very feature, DNN generally reconstruct the invariant mass even if it is not directly included in training variables. As a result, the distributions of categorized events are largely distorted. We used PyTorch Paszke et al. (2019) to implement DNN and used SCIKIT LEARN Pedregosa et al. (2011) for preprocessing. Selected models are given in Appendix.B.2. Fig.1 shows the actual distortion by ordinary DNN discriminator. In categories with high signal purity, even the background distribution resembles the resonance peak of the signal. It can mimic signals and overestimate the number of signal events. Therefore, the extraction of the number of events by fitting to $m_{\mu\mu}$ is not valid any more under such distortion. In such cases, it would be more appropriate to extract the number of events by fitting signal strength modifier to the distribution of output directly. However, this method suffers from large systematic uncertainties from simulation and training and therefore fittint to the invariant mass is more preferred generally.
(a) DNN; without jet substructure variables($S_{n_{\text{jet}}}$)
(b) DNN; with jet substructure variables($J_{n_{\text{jet}}}$)
Figure 1: $m_{\mu\mu}$ distribution binned by optimal boundaries chosen by output of (ordinary) DNN discriminators.
@@
The CMS collaboration used boosted decision tree (BDT) to categorize their ggH data. The tree-based models including BDT are essentially complex collections of cuts on training variables. Therefore, generally BDT has good performance for theoretically well-motivated training variables such as $p_{T},\eta$ and so on. It is also easy to interpret the result. However, compared to network-based models, it lacks the ability of capturing complex non-linear relationships among training variables. Generally it is the reason why network based discriminators such DNN are preferred for more complex data. Ironically, this intrinsic weakness of tree-based model would be profitable for making the resultant discriminator to be independent to $m_{\mu\mu}$.
To compare our result, we also conducted the BDT discrimination. We used the gradient-boosted tree implemented in SCIKIT LEARN. The Friedman MSE is used as a splitting criterion and a deviance is used as a loss. Selected models are given in Appendix.B.1. Fig.2 shows $m_{\mu\mu}$ distribution categorized by output of BDT discriminators. It is obvious that the $m_{\mu\mu}$ distributions are much more similar compared to the ordinary DNN case(Fig.1). However, the distributions are not exactly the same. There is a difference in slope of background distributions which also can be problematic for extracting the number of signal events. Actually the CMS collaboration was aware of it and use mass side band events to constrain the background distribution with modified fitting function having common shape and category specific slope (The ‘core-pdf’ method).
(a) BDT; without jet substructure variables($S_{n_{\text{jet}}}$)
(b) BDT; with jet substructure variables($S_{0},J_{1}$ and $J_{2}$)
@@
Figure 2: $m_{\mu\mu}$ distribution binned by optimal boundaries chosen by output of BDT discriminators.
For both BDT and DNN, the discrimination result is not statistically independent to $m_{\mu\mu}$. To achieve the independence, we exploit the flexibility of DNN. The network in DNN is an analytic function and it can handle wide variety of loss functions, which only requires to be differentiable. Therefore we included ‘decorrelation loss’ $\mathcal{L}_{\text{decor}}$ with a parameter $\lambda$ as $\mathcal{L}_{\text{tot}}\equiv\mathcal{L}_{\text{discrim}}+\lambda\mathcal{L}_{\text{decor}}$, where the discrimination loss $\mathcal{L}_{\text{discrim}}$ is the binary cross entropy. The $\mathcal{L}_{\text{discrim}}$ and $\mathcal{L}_{\text{decor}}$ are in conflict since $m_{\mu\mu}$ has strong discrimination power. However what we want to achieve is the decorrelation even if it costs discrimination performance. It can be achieved by setting $\lambda$ to be large enough. The distance correlation ($\equiv\mathcal{R}$) is used as a measure of statistical dependence. We required the discrimination score to be independent to $m_{\mu\mu}$ separately for signal and background events, $\mathcal{L}_{\text{decor}}\equiv\mathcal{R}(m_{\mu\mu},\text{score}|\text{signal})+\mathcal{R}(m_{\mu\mu},\text{score}|\text{background})$. We will call this DNN model with decorrelation loss as ‘decorrleated DNN’.
The decorrelated DNN is implemented with PyTorch again. We added decorrelation loss term every two epochs for calculation efficiency. Furthermore, since the distance matrix requires large memory, 1% of every batch is sampled to calculate $\mathcal{R}$. To test the degree of decorrelation at the end of training, the distance correlation is calculated with sample of size 25,000 for signal and background respectively (which is actually the limitation of our gpu memory). Selected models are given in Appendix.B.3. Fig.3 shows the $m_{\mu\mu}$ distributions which are not distorted by the categorization. Therefore, there will be no systematic error from dependence.
(a) decorrelated DNN; without jet substructure variables ($S_{n_{\text{jet}}}$)
(b) decorrelated DNN; with jet substructure variables ($J_{n_{\text{jet}}}$)
@@
Figure 3: $m_{\mu\mu}$ distribution binned by optimal boundaries chosen by output of decorrelated DNN discriminators.

### v2

@@ Abstract
Discovery of the Higgs boson decay to dimuon is anticipated soon based on the current evidence. Precise categorization of events without affecting the invariant mass shape is crucial in the analysis. Decorrelation of the invariant mass and the output of discriminators (the score of discriminators) is essential for consistent and precise analysis. In this paper, we use distance correlation as the additional loss function to achieve the decorrelation for discriminators and examine various analysis methods. The significance of the Higgs to dimuon signal from gluon fusion is considerably improved by including jet substructure variables.
@@ I Introduction
After Higgs discovery Chatrchyan et al. (2012); Aad et al. (2012), various properties of Higgs have been measured including Higgs couplings to Standard Model (SM) particles. The SM predicts the Higgs coupling to a particle is proportional to its mass. Accordingly, Higgs couplings to heavy particles such as gauge bosons and third generation fermions are well measured Aad et al. (2020); Aaboud et al. (2018); Sirunyan et al. (2018); Aaboud et al. (2019). However, couplings to relatively light particles have not been satisfactorily measured so far. The Higgs to dimuon channel is the most sensitive channel among the second generation fermions at the LHC due to the relatively clean leptonic final states. The property of Higgs couplings to light fermions can verify the Higgs mechanism of the SM more accurately. Furthermore, the growing interest towards a muon collider makes this channel more interesting.
@@
A crucial requirement for the categorized invariant mass fitting is that the information about the invariant mass should not be used in the categorization process. In other words, the output of discriminators, or the discriminator score, should be statistically independent of the invariant mass. Categorization by a score which is statistically dependent on the invariant mass generally results in different invariant mass distributions for each category, $f(m|\text{score})\neq f(m)$, where $f$ represents the probability density function. If the invariant mass distribution is distorted, the number of events extracted from each category can not be reliably trusted. In extreme cases, a resonance peak near the Higgs mass may appear in the background $m_{\mu\mu}$ distribution, adding significant uncertainties on the number of signal events near the Higgs bump region. In addition, if we fit to the invariant mass after the categorization by a discriminator correlated to the invariant mass, it results in double usage of the invariant mass, compounding the error. Therefore, the standard likelihood fitting to the invariant mass distribution can be consistently combined with machine discriminator results only when two results are decorrelated/deassociated.  [^1] Technically, the term ‘deassociate/deassociation’ is more appropriate Altman and Krzywinski (2015), but we will use the term ‘decorrelate/decorrelation’ following other literature
@@
To avoid the distortions in the invariant mass distribution that arise when events are selected by a discriminator correlated with the invariant mass, researchers have not only securely excluded the invariant mass, but also carefully selected other variables to make the categories as insensitive to the invariant mass as possible. However, as will be seen in the paper, machine discriminators generally learn about invariant mass since it has a strong discrimination power. An ad hoc choice of variables which seemingly unrelated to the invariant mass, qualitative check of distortion in distributions after categorization Bourilkov et al. (2019) and a posteriori modification of fitting functions are not satisfactory solutions. In this paper we provide a quantitative and consistent decorrelation framework to deal with this problem.
There are numerous techniques of decorrelation have been developed to decouple a discriminator from a specific feature Louppe et al. (2016); Shimmin et al. (2017); Englert et al. (2019); Clavijo et al. (2021); Klein and Golling (2022); Algren et al. (2024); Dolen et al. (2016); Moult et al. (2018); Windischhofer et al. (2020); Kitouni et al. (2020); Mikuni et al. (2022); Xia (2019); Kasieczka et al. (2021); Collaboration (2018, 2020); Wunsch et al. (2020). Among those various methods, we choose to use a measure of dependence ‘distance correlation’ first used for decorrelation in High Energy Physics by Kasieczka and Shih (2020). The distance correlation is capable of capturing both linear and non-linear dependence, well-defined in both population and sample sapces, and differentiable, so being compatible with deep learning frameworks. At the same time, the range is well defined such that it is 0 when there is no correlation and is 1 when it is fully deterministic/dependent. By leveraging the flexiblility of deep neural network (DNN), we use the distance correlation between $m_{\mu\mu}$ and output of discriminator as an additional loss for training. As a result, it has been found that almost perfect statistical independence between $m_{\mu\mu}$ and the output of discriminator can be systematically achieved for both signal and background data classes simultaneously. It allows us to choose the required degrees of independence, which is crucial for accurate analysis, despite being in a trade-off relation to the performance of the discriminator. This approach ensures reliable categorization without compromising the integrity of the invariant mass distribution.
@@
Compared to the analysis using only the kinematic variables of reconstructed particle objects, incorporating jet substructure variables of the jets associated with the dimuon system as additional input variables can provide extra power, especially in discriminating the gluon dominant gluon fusion (ggH) signal from other background processes. In Cho et al. (2020), it has been demonstrated that the ggH channel can be exploited to improve the sensitivity of the Higgs invisible signal if jet substructure variables are fully utilized to distinguish the flavors and kinematics of initial state radiation (ISR) by DNN analysis. The differences in parton luminosities and QCD color factors of vertices make the ggH ISR to be gluon-rich, while make the background Drell-Yan (DY) ISR to be quark-rich. In addition, due to the helicity conservation, ggH ISR has a deficit of quark jets at the central rapidity region [^2] There is no issue of decorrelation in Higgs to invisible study. The treatment of systematic uncertainties originating from the simulation remain as an open question.. In the Higgs to dimuon discovery, currently the vector boson fusion (VBF) is the leading production channel despite the fact that the ggH production cross section is more than 10 times larger than VBF one at $\sqrt{s}=13$TeV CERN (2017). Therefore we included the jet substructure variables to improve the sensitivity of Higgs to dimuon signal from ggH channel and compared the results.
@@
The paper is composed as following. In section II, we first outline the event simulation and ggH selection cuts. We also summarize the analysis strategy including training variables, common training schemes, and the categorization procedure. In section III, we introduce three types of discriminators and compare their result in detail. In particular, with the $m_{\mu\mu}$ decorrelated discriminator implemented using the distance correlation, the degree of decorrelation and the discrimination performance will be compared. In section IV, we compute the discovery significance using the categorized events and check the effect of including jet substructure variables.
@@ II Events and analysis strategy
Event simulation
When an event is classified as being from ggH, it means that the event passed all selection cuts to identify ggH, regardless of its genuine origin. Therefore, all possible Higgs production channels should be simulated. However, the contribution of other channels to the identified ggH signals is much smaller than that of the genuine ggH channel, as their cross sections are much smaller. The VBF channel has the next-to-leading-order contribution which is less than 5% of the ggH channel. Thus we simulated only the ggH and VBF channels. Similarly, there are many channels that contribute as backgrounds to ggH. Among them, we included DY, which is the dominant background and $\bar{\text{t}}$, which is more than 20% of DY for multi-jet events. All the events were simulated by MadGraph5_aMC@NLO Alwall et al. (2014) at $\sqrt{s}=13$ TeV and NLO in QCD. Each simulation was done inclusively for the number of jets, so that we can properly include both soft and hard jet emissions.
Event selection
After the simulation, events are selected for ggH identification by the criteria as the following. Basically, an event is required to have two muons with opposite charges, each with $p_{T}>20$GeV and $|\eta|<2.4$. Any events including b-jet and/or additional charged lepton are vetoed to minimize overlap with $\bar{t}$ and VH channels. Finally, events with two energetic jets with a large separation in pseudorapidity ($m_{jj}>400$GeV, $|\Delta\eta_{jj}|>2.5)$ are rejected as these are more likely to be identified as VBF.
Training variables
To improve the signal-to-background ratio, various information other than $m_{\mu\mu}$ is used to divide the total data into some categories. Recently, with the development of machine learning techniques, output of machine learning discriminators is frequently used for categorization and allows us to use multiple variables simultaneously.
The training variables used by the CMS collaboration are denoted by $S_{n_{\operatorname*{\text{jet}}}}$ for each jet multiplicity $n_{\operatorname*{\text{jet}}}$. $S_{n_{\text{jet}}}$ are carefully selected to exclude $m_{\mu\mu}$ information. See Appendix A.1 for more details. Please note that $n_{\text{jet}}=2$ means events with at least two jets, and the kinematics of jets are included in terms of pseudorapidity $\eta$ of leading and next-to-leading jets.
@@
Training and categorization
The discriminators are trained separately for different $n_{\operatorname*{\text{jet}}}$. One million signal and background events each are used for training (70%) and validation (30%). Three different machine discriminators are trained for comparison.
After training, total data is categorized based on the output of the discriminators, or equivalently, by the signal efficiency. Boundaries are recursively chosen to maximize total expected significance ($S/\sqrt{B}$) as a figure of merit, while minimizing the number of categories until the gain from an additional boundary is less than 1%.
Absence of shape fitting
Typically, the next step involves fitting data to the $m_{\mu\mu}$ distribution to extract the number of events. However we will skip the extraction procedure and use the number of events by simulation directly because we do not have real observed data available for fitting. Fitting simulated data again would not be worth much since we already used large dataset. While our results rely on simulation, it is sufficient for comparing discrimination models with and without decorrelation, and assessing the impact of jet substructure variables featuring the properties of the jets associated with the dimuon system. Therefore after the categorization, we will compute the discovery significance by profile likelihood fitting to the $m_{\mu\mu}$ distribution.
@@ III Discriminators
Boosted Decision Tree
@@
The CMS collaboration used boosted decision tree (BDT) to categorize their ggH data. A tree-based model such as BDT is essentially a complex collection of cuts on training variables. This makes BDTs generally well-suited for tasks with theoretically well-motivated variables such as transverse momentum ($p_{T}$), pseudo rapidity ($\eta$) and so on. It is also easy to interpret. However, compared to network-based models like DNN, it lacks the ability of capturing complex non-linear relationships among training variables. Generally it is the reason why network based discriminators are preferred for more complex data analysis. Ironically, this intrinsic weakness of tree-based model would be profitable for making the resultant discriminator to be independent to $m_{\mu\mu}$.
For comparison, we also performed BDT discrimination using the gradient-boosted tree provided by SCIKIT LEARN with Friedman mean squared error as the splitting criterion and deviance as the loss function. The details of trained model can be checked in Appendix C.1. Fig. 1 shows $m_{\mu\mu}$ distributions categorized by the output of BDT discriminators.
@@
Figure 1: BDT; Normalized $m_{\mu\mu}$ distributions binned by optimal boundaries chosen by output of BDT discriminators trained without jet substructure variables ($S_{n_{\text{jet}}}$).
The $m_{\mu\mu}$ distributions categorized by BDT are similar across the categories, implies that the choice of input variables having no or weak correlation with $m_{\mu\mu}$ worked relatively well for BDT. However, the non-zero correlation manifests itself in different slopes of background distributions. It is problematic for extracting the number of signal events. This issue was already recognized by the CMS collaboration. They used mass sideband events to constrain the background distribution with a modified fitting function having common shape and category specific slope, known as the ‘core-pdf’ method. We tried to eliminate the correlation quantitatively by including the distance correlation loss without a posteriori prescription. The definition of distance correlation is given in Appendix B. Unfortunately, the distance correlation is incompatible with BDT. The BDT requires decomposability of loss, but the distance correlation can not be defined locally for each point. On the other hand, the DNN can accommodate it, so we conducted DNN with and without decorrelation loss too.
Deep Neural Network
The astonishing success of DNN is based on the ability to learn powerful hidden representations from training variables. However, due to the very feature, DNN can reconstruct the invariant mass even if it is not directly included in training variables. As a result, the distributions of categorized events can be largely distorted.
We used PyTorch Paszke et al. (2019) to implement DNN and used SCIKIT LEARN Pedregosa et al. (2011) for preprocessing the training data. The details of trained model can be checked in Appendix C.2. Fig. 2 shows the actual distortion by ordinary DNN discriminator. In categories with high signal purity, even the background distribution resembles the resonance peak, which can mimic signals and overestimate the signal counts. The severe distortion in DNN result implies that the DNN is capable of reconstructing $m_{\mu\mu}$ from input variables which have complex non-linear relationships with $m_{\mu\mu}$. It is not reliable to extract the number of events by fitting to $m_{\mu\mu}$ under such distortion.
@@
Figure 2: DNN; Normalized $m_{\mu\mu}$ distributions binned by optimal boundaries chosen by output of (ordinary) DNN discriminators trained without jet substructure variables ($S_{n_{\text{jet}}}$).
In such cases, fitting a signal strength modifier directly to the output of discriminator (score) distribution might be a more tailored approach, but this approach often comes with significant drawbacks. This method is susceptible to larger systematic uncertainties that originate from two sources: the inherent uncertainties in Monte Carlo simulation and the uncertainties from training algorithms themselves. These uncertainties lead to a large variance in the resulting score distribution and make it difficult to extract reliable signal strengths.
Consequently, fitting to the invariant mass distribution remains to be the preferred method, being simpler and more robust in most cases. As a single well-understood variable, the invariant mass is less prone to the uncertainties that distort the score distribution, and this makes the invariant mass fit a more reliable indicator for extracting signal strengths and further analysis. Therefore, a proper implementation of decorrelation is inevitable to achieve high precision.
Decorrelated DNN
For both BDT and DNN, the discrimination result is not statistically independent to $m_{\mu\mu}$. To achieve the independence, we exploit the flexibility of DNN. The network in DNN is an analytic function and it can handle wide variety of loss functions, which only requires to be differentiable. Therefore we introduce a ‘decorrelation loss’ $\mathcal{L}_{\text{decor}}$ with a parameter $\lambda$ as $\mathcal{L}_{\text{tot}}\equiv\mathcal{L}_{\text{discrim}}+\lambda\mathcal{L}_{\text{decor}}$, where the discrimination loss $\mathcal{L}_{\text{discrim}}$ is the binary cross entropy. Here, $\mathcal{L}_{\text{discrim}}$ and $\mathcal{L}_{\text{decor}}$ are inherently conflicting. While $\mathcal{L}_{\text{discrim}}$ prioritizes distinguishing signal from background, $\mathcal{L}_{\text{decor}}$ aims to weaken the correlation between the discrimination score and $m_{\mu\mu}$. To achieve such decorrelation, we can increase the weight of $\mathcal{L}_{\text{decor}}$ by setting a sufficiently large $\lambda$ value, at the expense of some discrimination performance.
The distance correlation ($\equiv\mathcal{R}$) is used as a measure of statistical dependence. We required the discrimination score to be independent to $m_{\mu\mu}$ separately for signal and background events, $\mathcal{L}_{\text{decor}}\equiv\mathcal{R}(m_{\mu\mu},\text{score}|\text{signal})+\mathcal{R}(m_{\mu\mu},\text{score}|\text{background})$. We will call this DNN model with decorrelation loss as ‘decorrleated DNN’ from now on.
The decorrelated DNN is implemented with PyTorch again. We applied the decorrelation loss term every two epochs for calculation efficiency. Furthermore, since the distance matrix for the computation of distance correlation requires large memory, 1% of every batch is randomly sampled to compute $\mathcal{R}$. To test the degree of decorrelation at the end of training, the distance correlation is calculated with sample of size 25,000 for signal and background respectively due to GPU memory limitations. Selected models are given in Appendix C.3. As shown in Fig. 3, the $m_{\mu\mu}$ distributions for the decorrelated DNN exhibit minimal distortion across categories. This indicates a successful reduction in the dependence between the discrimination score and the invariant mass. Consequently, the systematic errors arising from this dependence can be significantly mitigated.
Figure 3: decorrelated DNN; Normalized $m_{\mu\mu}$ distributions binned by optimal boundaries chosen by output of decorrelated DNN discriminators trained without jet substructure variables ($S_{n_{\text{jet}}}$).